suppose the lengths of the pregnancies' of a certain animal are approx. normally distributed with a mean = 255 days and standard deviation of 14 days

what is the probability that a random sample of 48 pregnancies has a mean gestation of 250 days or less?

(d) the probability that the mean of a random sample of 48 pregnancies is less than 250 days is approx. ___?
Answer #1

Solution :

Given that ,

mean = \mu = 255

standard deviation = \sigma = 14

\sigma\bar x = \sigma / \sqrt n = 14 / \sqrt 48 = 2.0207

P(\bar x\leq 250) = P((\bar x - \mu \bar x ) / \sigma \bar x\leq(250 - 255) / 2.0207)

= P(z \leq -2.47)

= 0.9932

Probability = 0.9932

(d)

P(\bar x < 250) = P((\bar x - \mu \bar x ) / \sigma \bar x < (250 - 255) / 2.0207)

= P(z < -2.47)

= 0.9932

The probability that the mean of a random sample of 48 pregnancies is less than 250 days is approx.0.9932